Mike kicks soccer ball at a speed of 2.5m/s his friend Jake kicks it back at a speed of 5.5 m/s (Mass of ball is 0.5kg) a.What was the change in momentum? b. If the ball was in contact with Jakes foot for 0.07 seconds, what force did Jake apply?


1
Expert's answer
2021-05-04T06:34:07-0400

The information given is

  • The initial velocity is "\vec{V_{i}}=+2.5\;\text{m}/\text{s}"
  • The final velocity "\vec{V_{f}}=-5.5\;\text{m}/\text{s}" city is
  • The mass is "m=0.5\;\text{kg}"
  • The interaction time is "t=0.07\;\text{s}"

Part A


The change in momentum is given by


"\Delta \vec{P}=\vec{V}_{f}\;m-\vec{V}_{i}\;m"


Where.

"V_{i}" is the initial velocity

"V_{f}" is the final velocity.

"m" is the mass.


Evaluating numerically.


"\Delta \vec{P}=\vec{V}_{f}\;m-\vec{V}_{i}\;m\\"

"\Delta \vec{P}=-5.5\;\text{m}/\text{s}_{f}\times 0.5\;\text{Kg}-2.5\;\text{m}/\text{s}_{f}\times 0.5\;\text{Kg}\;m\\"

"\Delta \vec{P}=-4.0\;\text{Kg}\;\text{m}/\text{s}"


Answer A

The change in momentum is "\displaystyle \color{red}{\boxed{\Delta P=-4.0\;\text{Kg}\;\text{m}/\text{s}}}"


Part B


The average force is


"\vec{F}_{avg}=\dfrac{\Delta \vec{P}}{t}"


Where.

"\Delta \vec{P}" is the change of momentum.

"t" is the interaction time.


Evaluating numerically.


"\vec{F}_{avg}=\dfrac{\Delta \vec{P}}{t}\\
\vec{F}_{avg}=\dfrac{-4.0\;\text{Kg}\;\text{m}/\text{s} }{0.07\;\text{s}}\\
\vec{F}_{avg}=-57\;\text{N}"


Answer B

The average force is "\displaystyle \color{red}{\boxed{\vec{F}_{avg}=-57\;\text{N}}}"


The negative sign indicates that the force and the change in momentum (impulse) points in the opposite direction to the initial movement, that is, when it goes back towards Mike